# Negative Preference Optimization: From Catastrophic Collapse to Effective Unlearning

### v1

@@ 1 Introduction
There is a vast literature on machine unlearning and LLM unlearning. Due to limited space, we discuss these related work in Appendix 1.1.
@@ 1.1 Related work
Since its proposal by Cao & Yang (2015), machine unlearning has been extensively studied in the classification literature (Bourtoule et al., 2021; Golatkar et al., 2020; Ginart et al., 2019; Thudi et al., 2022; Izzo et al., 2021; Koh & Liang, 2017; Guo et al., 2019; Sekhari et al., 2021). For reviews of existing works, see Liu et al. (2024a); Zhang et al. (2023); Nguyen et al. (2022); Xu et al. (2023); Si et al. (2023). In particular, Ginart et al. (2019); Guo et al. (2019); Sekhari et al. (2021) introduced theoretical metrics for machine unlearning based on the notion of differential privacy and proposed provably efficient unlearning methods based on Newton update removal mechanisms. However, these algorithms require computing the Hessian of loss functions, which is intractable for LLMs.
@@ 2.1 Catastrophic collapse of gradient ascent
Figure 2: Comparison between GA and NPO on forget quality, model utility, KL divergence on the real-world Set, and the answers to the forget set. The rightmost figure shows the answers generated from variants of GA and NPO that incorporates the RT loss. All figures are generated on the Forget05 task in the TOFU data, trained for 10 epochs (detailed setup in Section E.1).
@@ 3 Negative Preference Optimization
Figure 3: Retain difference versus forget difference for ${\mathrm{GA}}$ and ${\mathrm{NPO}}$ with varying levels of ${\beta}$ in the binary classification experiment with $\alpha=1$. The Pareto curves all start from the bottom right corner $(1.70,0.02)$ and are computed by averaging over $5$ instances. We observe that the ${\mathrm{NPO}}$ trajectory converges to the ${\mathrm{GA}}$ trajectory as ${\beta}\to 0.$ Here retain distance and forget distance denote the KL divergence between the distributions of the predictions of the retrained and the unlearned model, on the retain and the forget distribution, respectively. More details can be found in Section 4.
@@
$\nabla_{\theta}\mathcal{L}_{{\mathrm{GA}}}=$ $~{}-\mathbb{E}_{{{\mathcal{D}}_{\rm FG}}}[\nabla_{\theta}\log{\pi}_{\theta}({y}|{x})],$ (4)
$\nabla_{\theta}\mathcal{L}_{{\mathrm{NPO}},{\beta}}=$ $~{}-\mathbb{E}_{{{\mathcal{D}}_{\rm FG}}}[\mathsf{W}_{\theta}({x},{y})\nabla_{\theta}\log{\pi}_{\theta}({y}|{x})],$ (5)
@@ 5.1 Experimental setup
Dataset and Metric.
We evaluate unlearning methods on the TOFU dataset (Maini et al., 2024). It contains 200 fictitious author profiles, each consisting of 20 question-answer pairs. TOFU introduces three levels of tasks, each aiming to forget 1% , 5% , and 10% of the data, referred to as Forget01, Forget05, and Forget10, respectively. We measure the effectiveness of unlearning methods via Forget Quality and Model Utility as in Maini et al. (2024). Forget quality assesses how well the unlearned model mimics the retrained model (defined as the model trained only on the retain set), while model utility measures the general capacities and the real-world knowledge of the unlearned model. Since the forget quality is defined as the p-value of the Kolmogorov-Smirnov test, which tests the similarity between some distributions generated by the unlearned model and the retrained one, we treat a forget quality greater than $0.05$ as evidence of a meaningful forgetting. More details are deferred to Section E.1.1 and Section E.1.2.
@@
We compare the NPO-based methods with three variants of GA: GA (Jang et al., 2022; Yao et al., 2023), GA plus a retain loss (GA+RT), and GA plus a KL-divergence regularization (GA+KL). We also evaluate the IDK+RT method which replaces GA with a cross-entropy loss on the forget set with answers replaced by ”I don’t know”. Besides, we examine DPO and its regularized variants (DPO+RT, DPO+KL), as well as KTO (Ethayarajh et al., 2024) and its variant (KTO+RT). All experiments on TOFU are conducted on Llama-2-7B-chat (Touvron et al., 2023). See Section E.1 for more details.
@@ 5.2 Results
Figure 5 illustrates the trade-off between forget quality and model utility for various unlearning methods in the Forget01, Forget05, and Forget10. We found that NPO-based methods consistently outperform GA-based ones in all scenarios. Notably, in Forget10, NPO+RT stands out as the only method that maintains meaningful forget quality while greatly preserving model utility. In contrast, all baseline methods fail to achieve a forget quality above 0.05.
@@
NPO avoids catastrophic collapse.
Figure 6 illustrates the evolution of forget quality and model utility along the unlearning process. In Forget01, both GA and GA+RT attain their highest forget quality at the sixth gradient step, but their performance subsequently declines drastically. Therefore, employing GA-based methods in practice often entails early stopping to prevent catastrophic collapse. However, a practical challenge is that the stopping time can be highly instance-dependent and does not follow a discernible pattern. In contrast, NPO-based methods display considerably greater stability, with forget quality consistently reaching and maintaining a plateau.
@@
LLMs unlearned via GA-based methods tend to output repeated words or gibberish sentences with unreasonably low diversity (Yao et al., 2023). Moreover, IDK and DPO-based methods tend to show excessive ignorance. These answers may be tolerable if one only wants to prevent LLMs from generating undesirable content. Still, they will definitely be unsatisfactory under the stronger goal of approximate unlearning, which aims to mimic the retrained model. We show in Figure 7 that NPO+RT outputs incorrect sentences with similar templates for questions in the forget set while generating fluent and correct answers for other questions, greatly enhancing the fluency and diversity of the generated content.
@@
To further investigate the role of retain loss beyond Maini et al. (2024), we evaluate NPO+RT with the weights of the retain loss varying from 0 to 5 (Figure 11). While it is natural that adding retain loss improves the model utility, we are surprised that the forget quality also grows. Specifically, the forget quality increases as the weight of the retain loss grows from 0 to 2. We conjecture that the retain loss term helps the model preserve answer templates and linguistic structures, while the NPO term forces the model to forget some specific facts. Combining these two effects pushes the model to approximate the retrained model by generating outputs with similar templates but incorrect entities. We also note that further increasing the weight of the retain loss (e.g., from 2 to 5) leads to a drop in forget quality.
@@ 5.3 Forgetting beyond 10% of TOFU
Having demonstrated that NPO-based methods can effectively unlearn 10% of the TOFU data, we now expand our scope to the tasks of forgetting 20%, 30%, and 50% of the TOFU data (referred to as Forget20, Forget30, Forget50, respectively). Details about the extended dataset are deferred to Section E.1.1. We show in Section E.2 that NPO+RT is the sole method to exhibit meaningful forget quality (a p-value above 0.05) in Forget20 and Forget30. Even in Forget50, where the vanilla NPO+RT achieves a forget quality around $10^{-3}$, it still significantly outperforms other methods.
@@ Acknowledgement
Song Mei is supported by NSF DMS-2210827, CCF-2315725, NSF Career DMS-2339904, and a Google Research Scholar Award. The authors would like to thank Baihe Huang, Xuelin Yang for the valuable discussions. The authors would like to thank Jiantao Jiao for sharing his GPU resources. This research was supported by the Center for AI Safety Compute Cluster. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.
@@ B.1 Synthetic Experiment
Figure 10: The evolution of the Forget KL during the unlearning process on the Forget10 task in TOFU data. Note that the KL term in GA+KL is the divergence on the retain set, not the forget set. More experimental details are included in Section E.1.
B.2 Experiments on the TOFU dataset
We also examine the Forget KL during the unlearning process in the TOFU dataset. We first observed that while GA and GA+RT tend to induce an explosively large KL divergence on the forget set during the unlearning process, the NPO-based approaches induce a much slower growth of Forget KL (Figure 10). It stabilizes at a moderate level even after several epochs. One natural insight from this distinction is that even in the context of unlearning, a larger KL divergence on the forget set is not necessarily advantageous. Rather, a moderate and stabilized KL divergence on the forget set is preferable, which ensures the unlearned models generate fluent outputs with reasonable linguistic structures but incorrect content. This also suggests that KL divergence on the forget set may not be a suitable objective function to maximize for unlearning LLMs, contrary to what was done in some prior literature (Chen & Yang, 2023).
Appendix C Ablation experiments on the role of retain loss
Figure 11: The evolution of the forget quality and model utility when we adjust the weights of the NPO term and the retain loss term in NPO+RT. The experiments are performed on Forget10.
Maini et al. (2024) demonstrated that methods incorporating a retain set outperform those that solely optimize a loss function based on the forget set. It is natural that a regularization via retain loss will improve the performance on the retain set, which agrees with the tendency shown in Figure 11 that the model utility monotonically grows as the weight of the retain loss increases. Surprisingly, with the help of the retain loss, the forget quality can also improve. Specifically, NPO+RT shows a much more significant forget quality as the weight of the retain loss grows from 0 to 2 (see the top-right panel of Figure 6 and Figure 11). We conjecture that it is because fine-tuning the retain loss term can help the model to preserve answer templates and their linguistic structure while optimizing the NPO term forces the model to forget some specific facts. The combination of the two pushes the unlearned model to approximate the retrained model by generating the outputs with similar templates but incorrect entities. However, further increasing the weight of the retain loss (e.g., from 2 to 5, in Figure 11) leads to a drop in the forget quality, possibly due to the diminished scale of the NPO term. Notably, in our experiments, the retain loss plays a more significant role when we target forgetting a larger fraction of the data (See the middle and right panels of Figure 6).
Appendix D Experimental details of the synthetic experiments
@@
Appendix E Experiments on the TOFU dataset
@@
E.1 Experiments Setup
E.1.1 Dataset
@@ E.1.3 Baseline Methods
E.1.4 Hyperparameters
For the experiments TOFU, we use Llama2-7b-chat model (Touvron et al., 2023). All experiments are conducted with two A100 GPUs. We use AdamW with a weight decay of $0.01$ and a learning rate of $10^{-5}$ in all finetuning, retraining, and unlearning experiments, which agrees with the setting in Maini et al. (2024). We use an effective batch size of $32$ for all experiments. In finetuning and retraining, we train for 5 epochs, while we train for 10 epochs in the unlearning process. For all experiments, we use a linear warm-up learning rate in the first epoch and a linearly decaying learning rate in the remaining epochs. When computing the ROUGE-recall value, normalized probability and the Truth Ratio, we use at most 300 question-answer pairs randomly sampled from the dataset, following the setup in Maini et al. (2024).
E.2 Full Results

### v2

@@ 1.1 Related work
There is a vast literature on machine unlearning and LLM unlearning. Since its proposal by Cao & Yang (2015), machine unlearning has been extensively studied in the classification literature (Bourtoule et al., 2021; Golatkar et al., 2020; Ginart et al., 2019; Thudi et al., 2022; Izzo et al., 2021; Koh & Liang, 2017; Guo et al., 2019; Sekhari et al., 2021). For reviews of existing works, see Liu et al. (2024a); Zhang et al. (2023); Nguyen et al. (2022); Xu et al. (2023); Si et al. (2023). In particular, Ginart et al. (2019); Guo et al. (2019); Sekhari et al. (2021) introduced theoretical metrics for machine unlearning based on the notion of differential privacy and proposed provably efficient unlearning methods based on Newton update removal mechanisms. However, these algorithms require computing the Hessian of loss functions, which is intractable for LLMs.
@@ 2.1 Catastrophic collapse of gradient ascent
Figure 2: Comparison between GA and NPO on forget quality, model utility, KL divergence on the real-world Set, and the answers to the forget set. The rightmost figure shows the answers generated from variants of GA and NPO that incorporates the RT loss. All figures are generated on the Forget05 task in the TOFU data, trained for 10 epochs (detailed setup in Section D.1).
@@ 3 Negative Preference Optimization
Figure 3: Retain distance versus forget distance for ${\mathrm{GA}}$ and ${\mathrm{NPO}}$ with varying levels of ${\beta}$ in the binary classification experiment with $\alpha=1$. The Pareto curves all start from the bottom right corner $(1.70,0.02)$ and are computed by averaging over $5$ instances. We observe that the ${\mathrm{NPO}}$ trajectory converges to the ${\mathrm{GA}}$ trajectory as ${\beta}\to 0.$ Here retain distance and forget distance denote the KL divergence between the distributions of the predictions of the retrained and the unlearned model, on the retain and the forget distribution, respectively. More details can be found in Section 4.
@@
$\nabla_{\theta}\mathcal{L}_{{\mathrm{GA}}}=$ $~{}\mathbb{E}_{{{\mathcal{D}}_{\rm FG}}}[\nabla_{\theta}\log{\pi}_{\theta}({y}|{x})],$ (4)
$\nabla_{\theta}\mathcal{L}_{{\mathrm{NPO}},{\beta}}=$ $~{}\mathbb{E}_{{{\mathcal{D}}_{\rm FG}}}[\mathsf{W}_{\theta}({x},{y})\nabla_{\theta}\log{\pi}_{\theta}({y}|{x})],$ (5)
@@ 5.1 Experimental setup
Dataset and metrics.
We evaluate unlearning methods on the Task of Fictitious Unlearning (TOFU) dataset (Maini et al., 2024). It contains 200 fictitious author profiles, each consisting of 20 question-answer pairs generated by GPT-4 based on some predefined attributes. These fictitious profiles do not exist in the pre-training data, providing a controlled environment for studying unlearning LLMs. TOFU introduces three levels of tasks, each aiming to forget 1% , 5% , and 10% of the data, referred to as Forget01, Forget05, and Forget10, respectively. We measure the effectiveness of unlearning methods via Forget Quality and Model Utility as in Maini et al. (2024). Forget quality assesses how well the unlearned model mimics the retrained model (defined as the model trained only on the retain set), while model utility measures the general capacities and the real-world knowledge of the unlearned model. Since the forget quality is defined as the p-value of the Kolmogorov-Smirnov test, which tests the similarity between some distributions generated by the unlearned model and the retrained one, we treat a forget quality greater than $0.05$ as evidence of a meaningful forgetting. More details are deferred to Section D.1.1 and Section D.1.2.
@@
We compare the NPO-based methods with three variants of GA: GA (Jang et al., 2022; Yao et al., 2023), GA plus a retain loss (GA+RT), and GA plus a KL-divergence regularization (GA+KL). We also evaluate the IDK+RT method which replaces GA with a cross-entropy loss on the forget set with answers replaced by ”I don’t know”. Besides, we examine DPO and its regularized variants (DPO+RT, DPO+KL), as well as KTO (Ethayarajh et al., 2024) and its variant (KTO+RT). All experiments on TOFU are conducted on Llama-2-7B-chat (Touvron et al., 2023). See Section D.1 for more details.
Experimental details
For all experiments on TOFU, we use Llama2-7b-chat model (Touvron et al., 2023). All experiments are conducted with two A100 GPUs. We use AdamW with a weight decay of $0.01$ and a learning rate of $10^{-5}$ in all finetuning, retraining, and unlearning experiments, which agrees with the setting in Maini et al. (2024). We use an effective batch size of $32$ for all experiments. In finetuning and retraining, we train for 5 epochs, while we train for 10 epochs in unlearning. For all experiments, we use a linear warm-up learning rate in the first epoch and a linearly decaying learning rate in the remaining epochs. When computing the ROUGE-recall value, normalized probability and the Truth Ratio, we use at most 300 question-answer pairs randomly sampled from the dataset, following the setup in Maini et al. (2024).
@@ 5.2 Results
Figure 5 illustrates the trade-off between forget quality and model utility for various unlearning methods in the Forget01, Forget05, and Forget10. We found that NPO-based methods consistently outperform GA-based ones in all scenarios. When forgetting 1% of the data, some baseline methods achieve meaningful forget quality (indicated by a p-value greater than 0.05). Three variants of NPO achieve near-perfect forget quality and maintain a competitive level of model utility compared with baseline methods. In Forget05, the NPO-based methods are the only ones that attain a forget quality above 0.05. Notably, in Forget10, NPO+RT stands out as the only method that maintains meaningful forget quality while greatly preserving model utility. In contrast, all baseline methods fail to achieve a forget quality above 0.05.
@@
NPO avoids catastrophic collapse.
Figure 6 illustrates the evolution of forget quality and model utility along the unlearning process. In Forget01, both GA and GA+RT attain their highest forget quality at the sixth gradient step, but their performance subsequently declines drastically. Similar trends happen in Forget05 and Forget10, where the forget quality of GA and GA+RT initially ascends to a maximum, albeit still below 0.05, before rapidly diminishing to an exponentially small magnitude. Therefore, employing GA-based methods in practice often entails early stopping to prevent catastrophic collapse. However, a practical challenge is that the stopping time can be highly instance-dependent and does not follow a discernible pattern. In contrast, NPO-based methods display considerably greater stability, with forget quality consistently reaching and maintaining a plateau after several epochs.
@@
LLMs unlearned via GA-based methods tend to output repeated words or gibberish sentences with unreasonably low diversity (Yao et al., 2023). Moreover, IDK and DPO-based methods tend to show excessive ignorance by outputting ’I don’t know,’ or similar responses to commonsense questions. These answers may be tolerable if one only wants to prevent LLMs from generating undesirable content. Still, they will definitely be unsatisfactory under the stronger goal of approximate unlearning, which aims to mimic the retrained model. We show in Figure 7 that NPO+RT outputs incorrect sentences with similar templates for questions in the forget set while generating fluent and correct answers for other questions, greatly enhancing the fluency and diversity of the generated content.
@@
Maini et al. (2024) demonstrated that methods incorporating a retain set outperform those that solely optimize a loss function based on the forget set. To further investigate the role of retain loss beyond Maini et al. (2024), we evaluate NPO+RT with the weights of the retain loss varying from 0 to 5 (Figure 8). While it is natural that adding retain loss improves the model utility, we are surprised that the forget quality also grows. Specifically, the forget quality increases as the weight of the retain loss grows from 0 to 2. We conjecture that the retain loss term helps the model preserve answer templates and linguistic structures, while the NPO term forces the model to forget some specific facts. Combining these two effects pushes the model to approximate the retrained model by generating outputs with similar templates but incorrect entities. However, further increasing the weight of the retain loss (e.g., from 2 to 5, in Figure 8) leads to a drop in the forget quality, possibly due to the diminished scale of the NPO term. Notably, in our experiments, the retain loss plays a more significant role when we target forgetting a larger fraction of the data (See the middle and right panels of Figure 6).
Figure 8: The evolution of the forget quality and model utility when we tune the weights of the NPO term and the retain loss term in NPO+RT. The experiments are performed on Forget10. We observe that when we increase the weight for the retain loss term, the model utility increases monotonically while the forget quality initially improves but then starts to deteriorate. We remark that altering the weights for loss components does not affect the effective learning rate since, in practice, AdamW is scale-invariant.
Forget KL: The larger, the better? ✖ ✖ ✖
We also examine the Forget KL during the unlearning process in the TOFU dataset. We first observed that while GA and GA+RT tend to induce an explosively large Forget KL along the unlearning process, the NPO-based approaches induce a much slower growth of Forget KL (Figure 9). It stabilizes at a moderate level even after several epochs. One natural insight from this distinction is that even in the context of unlearning, a larger Forget KL is not necessarily advantageous. Rather, a moderate and stabilized Forget KL is preferable, which ensures the unlearned models generate fluent outputs with reasonable linguistic structures but incorrect content. This also suggests that Forget KL may not be a suitable objective function to maximize for unlearning LLMs, contrary to what was done in some prior literature (Chen & Yang, 2023).
Figure 9: The evolution of the Forget KL during the unlearning process on the Forget10 task in TOFU data. Note that the KL term in GA+KL is the divergence on the retain set, not the forget set. More experimental details are included in Section D.1.
@@ 5.3 Forgetting beyond 10% of TOFU
Having demonstrated that NPO-based methods can effectively unlearn 10% of the TOFU data, we now expand our scope to the tasks of forgetting 20%, 30%, and 50% of the TOFU data (referred to as Forget20, Forget30, Forget50, respectively). Details about the extended dataset are deferred to Section D.1.1. We show in Section D.2 that NPO+RT is the sole method to exhibit meaningful forget quality (a p-value above 0.05) in Forget20 and Forget30. Even in Forget50, where the vanilla NPO+RT achieves a forget quality around $10^{-3}$, it still significantly outperforms other methods.
@@ Acknowledgement
Song Mei is supported by NSF DMS-2210827, CCF-2315725, NSF Career DMS-2339904, ONR N00014-24-S-B001, an Amazon Research Award, and a Google Research Scholar Award. The authors would like to thank Baihe Huang, Xuelin Yang for the valuable discussions. The authors would like to thank Jiantao Jiao for sharing his GPU resources. This research was supported by the Center for AI Safety Compute Cluster. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.
@@ B.1 Synthetic Experiment
Appendix C Experimental details of the synthetic experiments
@@
Appendix D Experiments on the TOFU dataset
@@
D.1 Experiments Setup
D.1.1 Dataset
@@ D.1.3 Baseline Methods
D.2 Full Results